Find the curvature of the space curve. r(t) = -5 i + (10 + 2t)j + (t? + 8) k Ov-2021 2052 or OK 2(+1312
Answer #1

Sol) F()x=0) + (1) Given, r(t) = -5i + (10+21); + (x2 +8)} = (-5, (10+21), (+2+8)) = r(t) = (0,2, 21) =1(t) = (0,0,2) Now,